A fuel contains 4% hydrogen; 89% carbon; 1% sulphur; 2,5% oxygen, 3,5% ash and 20% excess air is supplies

Calculate

1.1 The mass of oxygen required to complete combustion of 1kg of fuel
1
Expert's answer
2020-05-01T14:31:37-0400

1 kg of fuel contain 40 g of hydrogen, 890 g of carbon and 10 g of sulphur.


2H2 + O2 = 2H2O


For complete combustion of 40 g of hydrogen need 0.5*32*40/2= 320 g of oxygen.


C + O2 = CO2


For complete combustion of 890 g of carbon need 32*890/12 = 2373.3 g of oxygen.


S + O2 = SO2


For complete combustion of 10 g of carbon need 32*10/32 = 10 g of oxygen.


The mass of oxygen required to complete combustion of 1kg of fuel is 320+2373.3+10 = 2703.3 g.